OpenStudy (anonymous):

The value of the derivative of y=(x-1)(3x+2x^2) at x=1 is....

OpenStudy (anonymous):

Use product rule to find y'

OpenStudy (anonymous):

Do u know product rule.

OpenStudy (anonymous):

y = uv y' = u'v + uv'

OpenStudy (anonymous):

DOES IT HELP? @JosephLing

OpenStudy (anonymous):

still not understand

OpenStudy (anonymous):

Let u=x-1 AND v=(3x+2x^2)

OpenStudy (anonymous):

can u find u' and v'

OpenStudy (anonymous):

\[y'(x)=(3x+2x^2)+(x-1)(3+4x)=3x+2x^2+3x+4x^2-3-4x=6x^2+2x-3\\y'(1)=6(1)^2+2(1)-3=5\]

OpenStudy (campbell_st):

just expand it then differentiate y = 2x^3 + + x^2 - 3x the use the basic differentiate rule

OpenStudy (anonymous):

thx...everyone